OpenStudy (anonymous):

how to find the slope of (8,10), and (1,9)

hero (hero):

Using y = mx + b, insert the points one at a time into the formula: 9 = 1m + b 10 = 8m + b Isolate b in both equations b = 9 - m b = 10 - 8m set b = b 9 - m = 10 - 8m Place like terms on the same side: 8m - m = 10 - 9 Combine terms: 7m = 1 Isolate m: m = 1/7

OpenStudy (anonymous):

\[m=\frac{ (y2-y1) }{ (x2-x1) }\]

OpenStudy (anonymous):

What is the slope of the line passing through the points (2, 7) and (–1, 3)? Would the answer to this be 3/4?

OpenStudy (anonymous):

Almost... check to make sure you didn't reverse your x's and y's in the slope equation.

OpenStudy (anonymous):

 correct?